OpenStudy (anonymous):

Solve. 3x^2 – 7x – 2 = 0

OpenStudy (anonymous):

\[x=-5\pm \sqrt{13}/2 \]

OpenStudy (anonymous):

\[x=-3\pm \sqrt{13}\]

OpenStudy (anonymous):

\[1x=-5\pm \sqrt{37}/2\]

OpenStudy (anonymous):

\[x=5\pm \sqrt{13}/2\]

OpenStudy (anonymous):

these are the answers given to choose from

OpenStudy (anonymous):

x = [7 +- sqrt(49 - 4*3*-2)] / 6 = [ 7 +- sqrt73 ] / 2

OpenStudy (anonymous):

jimmy do u mean 37 cause they dont give an answer with sqrt 73

OpenStudy (anonymous):

mathmind can u help me with this please

OpenStudy (anonymous):

if you have an equation \[ax^2+bx+x=0\] use the quadratic equation to solve for x:\[Xroots= \frac{-b \pm \sqrt{b^2-4ac}}{2a} \] does that help ?

OpenStudy (anonymous):

sorry the first equation should be:\[ax^2+bx+c=0\]

jhonyy9 (jhonyy9):

is right the jimmyrep's answer just from last fraction the denominator is 6 and not is 2

OpenStudy (anonymous):

right jhony - the denominator is 2*3 = 6